OpenStudy (anonymous):

Find the center & radius for the following: 1. 2x^2+2y^2=72 2. (x-3)^2+(y-1)^2=4 3. (x-5)^2+(y+2)^2=49 4. (y-4)^2+(x-10)^2=64

OpenStudy (dumbcow):

these are all circles the circle equation is: (x-h)^2 +(y-k)^2 = r^2 center: (h,k) radius is r

OpenStudy (anonymous):

1)Center is (0,0) radius=6 2)Center is (3,1) radius=2 3)Center is (5,-2) radius=7 4)Center is (10,4) radius=8

OpenStudy (anonymous):

Your amazing! i have a few more.

OpenStudy (anonymous):

1. (x^2-6x+9)+(y^2+8y+16)=1

OpenStudy (anonymous):

Center is (3,-4) radius=1

OpenStudy (anonymous):

2. x^2+12x+y^2-8y=-27

OpenStudy (anonymous):

Center is (-6,4) radius=5

OpenStudy (anonymous):

Write an equation for the circle with center at (3,-4) and with (2,-1) a point on the circle.

OpenStudy (anonymous):

omg, amazinggggg. (:

OpenStudy (anonymous):

\[x^2+y^2-6x+8y+15=0\]

OpenStudy (anonymous):

My multiple choice answers are: (x-2)^2+(y+1)^2=1 (x+2)^2+(y-1)^2=10 (x-3)^2+(y+4)^2=10 (x-3)^2+(y+4)^2=5 (x+3)^2+(y-4)^2=100

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Your so smart.

OpenStudy (anonymous):

Glad to help